# On an entropy inequality for quadratic forms and applications

### v1

@@ Appendix A Appendix - Examples
In this appendix, we show details of some examples. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
Example A.1.
Given $0\leq s\leq 1$, and given $n$ sufficiently large, we will construct an $(s,O(1))$-Frostman random variable $X$ with $range(X)\subseteq[0,1)$, and compare $H_{n}(X)$, $H_{n}(X+Y)$, $H_{n}(X^{2}+Y^{2})$ and $H_{n}(XY)$, where $Y$ is an i.i.d. copy of $X$.
Let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. Take
@@
Denote by $\mathcal{L}^{d}$ the $d$-dimensional Lesbegue measure. Then $\mathcal{L}^{1}(A)=2^{-l}$. Let $X$ be a random variable having uniform distribution on $A$, i.e.,
@@
Claim 1. The random variable $X$ is $\left(s,O(1)\right)$-Frostman whenever $0\leq s\leq\frac{m}{l+m}$.
@@
when $l/m$ tends to infinity. For $l=0$ and $s=1$, we assert by Claim 2 that $H_{n}(X)=n=ns$. Combining Corollary 2.22, Lemma 2.36 and (11), we conclude that
@@
$H(X^{\prime}+Y^{\prime})\leq\log|range_{+}(X^{\prime}+Y^{\prime})|=m+O(1),$
@@
Recall that $X^{\prime}=D_{m+l}(X)$ and $Y^{\prime}=D_{m+l}(Y)$. Since the function $(x,y)\mapsto x^{2}+y^{2}$ on $[0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.19 that
@@
where $\mathcal{K}$ denotes the set consisting of integers $k$ such that $k=j_{1}^{2}+j_{2}^{2}$ for some $10\leq j_{1},j_{2}\leq 2^{m}-$ 1, and
@@
Since the function $(x,y)\mapsto xy$ on $[0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.19 that
@@
where $\mathcal{K}^{\prime}$ denotes the set consisting of integers $k$ such that $k=j_{1}j_{2}$ for some $10\leq j_{1},j_{2}\leq 2^{m}-$ 1, and
@@
Example A.2.
Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in A.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=\sqrt{X}$ and $\tilde{Y}=\sqrt{Y}$. By Example A.1, one sees that
@@
Example A.3.
Let $X$, $Y$ be the independent $(s,O(1))$-Frostman random variables defined in A.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=1+X$, which is $(s,O(1))$-Frostman with range contained in $[1,2)$. Let $\tilde{Y}$ be an i.i.d. copy of $X$. Take $\phi(x,y)=(x+y)^{2}$, which is a degenerate quadratic form. By Example A.1, one can deduce that
@@
and $Y$ has uniform distribution on $[0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is $(1/2,O(2^{m}))$-Frostman. We take $s=1/2$ and consider $n=2m$.
@@
Example A.5.
This example is an variant of Example A.4. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$. Note that $X$ and $Y$ are dependent. The marginal distributions of $X$ and $Y$ are both the uniform distribution on $\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$. So both $X$ and $Y$ are $(1/2,O(1))$-Frostman. Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$. We consider $n\leq 2m$.
@@
Moreover, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.32, if $21<s<3/$ 2, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with $s/2<t<s-1/2$. That is to say, the term $-\log C$ in Theorem 2.32 is essential.
@@
Example A.6.
@@
Let $X,Y$ be the random variables defined in Example A.1, and $X^{\prime}=D_{m+l}(X)$, $Y^{\prime}=D_{m+l}(Y)$, where the parameters are chosen as the following: for any given number $\eta>0$, let $m$ be any sufficiently large natural number, $l$ be a natural number with $(1+\eta/2)m\leq l\leq(1+\eta)m$, and $n=m+l$. By Example A.1, the random variables $X$ and $Y$ are $(s,O(1))$-Frostman, where $s=\frac{m}{l+m}\in[\frac{1}{2+\eta},1/2)$. Also, the equalities in (53) hold.
@@
Write $i=i_{1}-i_{2}$ and $i^{\prime}=i_{1}^{\prime}-i_{2}^{\prime}$. Then $|i|,|i^{\prime}|<2^{m}$ and $|i+\sqrt{2}i^{\prime}|<2^{-(1+\eta/2)m}$. Combining the fact that $(i,i^{\prime})\neq(0,0)$, one can deduce that neither $i$ nor $i^{\prime}$ can be 00.
@@
Appendix B Appendix - Existence of Certain Random Variable
@@
Lemma B.1.
Let $(A,B,C)$ be a given random variable. Let $h$ be a function such that $h(C)$ is a well-defined discrete random variable. Then there exists a random variable $(A^{\prime},B^{\prime},C^{\prime},B^{\prime\prime})$ such that $(A^{\prime},B^{\prime},C^{\prime})\equiv(A,B,C)$, $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))$, and $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$. Moreover, $B^{\prime\prime}$ and $B^{\prime}$ are conditionally i.i.d. given $h(C^{\prime})$. Furthermore, $(A^{\prime},B^{\prime})$ and $B^{\prime\prime}$ are conditionally independent given $h(C^{\prime})$.
@@
Suppose that $A,B,C$ are defined as measurable functions from a probability space $(\Omega,\mathcal{B},P)$ to some other measurable space. Now we construct a new probability space $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$. Let
$\Omega=\bigcup_{c\in range(h(C))}\Omega_{c},\quad\Omega_{c}:=\{h(C)=c\}=\{\omega\in\Omega:\,h(C(\omega))=c\}.$
@@
$\Omega^{\prime}=\bigcup_{c}(\Omega_{c}\times\Omega_{c}).$
Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{c}$ for some $c$, if and only if $h(C(\omega))=h(C(\nu))$. The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartisian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{c}\times\Omega_{c}$ for some $c$, then
$P^{\prime}(E\times F)=\begin{cases}\frac{P(E)P(F)}{P(\Omega_{c})},\quad&\text{if }P(\Omega_{c})>0,\\ 0,&\text{if }P(\Omega_{c})=0.\end{cases}$
In the following, we assume without loss of generality that $P(\Omega_{c})>0$ for all $c\in range(h(C))$. Now for any random variable $D$ defined on $(\Omega,\mathcal{B},P)$, let $D^{\prime}$ be the following random variable on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$:
$D^{\prime}(\omega,\nu)=D(\omega),\quad(\omega,\nu)\in\Omega^{\prime}.$
By this definition, one sees that $\big{(}f(D))^{\prime}=f(D^{\prime})$ for any admissible function $f$. In the following, we verify that $D^{\prime}\equiv D$. For any Borel set $L$,
$P^{\prime}(D^{\prime}\in L)$ $=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,D^{\prime}(\omega,\nu)\in L)$
$=\sum\limits_{c}P^{\prime}(\{(\omega,\nu)\in\Omega_{c}\times\Omega_{c}:\,D(\omega)\in L\})$
$=\sum\limits_{c}\frac{P(D\in L,h(C)=c)P(h(C)=c)}{P(h(C)=c)}=P(D\in L).$
In particular, we can take $D=(A,B,C)$ to get $(A^{\prime},B^{\prime},C^{\prime})\equiv(A,B,C)$. All their corresponding marginal distributions are also the same. One has $h(C)\equiv h(C^{\prime})$ as well.
Next, we define $B^{\prime\prime}$ as a function on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$, given by
$B^{\prime\prime}(\omega,\nu)=B(\nu),\quad(\omega,\nu)\in\Omega^{\prime}.$
@@
$P^{\prime}(B^{\prime\prime}\in L_{1},D^{\prime}\in L_{2},h(C^{\prime})=c)=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,B^{\prime\prime}(\omega,\nu)\in L_{1},D^{\prime}(\omega,\nu)\in L_{2},h(C^{\prime}(\omega,\nu))=c\})$
$\qquad=P^{\prime}(\{(\omega,\nu)\in\Omega_{c}\times\Omega_{c}:\,B(\nu)\in L_{1},D(\omega)\in L_{2}\})=\frac{P(D\in L_{2},h(C)=c)P(B\in L_{1},h(C)=c)}{P(h(C)=c)}.$
When we take $D$ to be trivial (a constant function), then $P^{\prime}(B^{\prime\prime}\in L_{1},h(C^{\prime})=c)=P(B\in L_{1},h(C)=c)$ and then $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))$. Moreover, the previous argument implies that $(D^{\prime},h(C^{\prime}))\equiv(D,h(C))$. We can calculate further to obtain
$P^{\prime}(B^{\prime\prime}\in L_{1},D^{\prime}\in L_{2},h(C^{\prime})=c)$ $=\frac{P(D\in L_{2},h(C)=c)P(B\in L_{1},h(C)=c)}{P(h(C)=c)}$
$=\frac{P^{\prime}(D^{\prime}\in L_{2},h(C^{\prime})=c)P^{\prime}(B^{\prime\prime}\in L_{1},h(C^{\prime})=c)}{P^{\prime}(h(C^{\prime})=c)}$
$=P^{\prime}(D^{\prime}\in L_{2},h(C^{\prime})=c)P^{\prime}(B^{\prime\prime}\in L_{1}|h(C^{\prime})=c).$
So $P^{\prime}(B^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(C^{\prime})=c)=P^{\prime}(B^{\prime\prime}\in L_{1}|h(C^{\prime})=c)$, i.e. $(B^{\prime\prime}|D^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$. Taking $D=(A,B)$, one gets $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$.
Moreover, if we take $D=B$, then $(B^{\prime\prime}|B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$. That is to say, $B^{\prime\prime}$ and $B^{\prime}$ are conditionally independent given $h(C^{\prime})$. Since $(B^{\prime\prime},h(C^{\prime}))\equiv(B,h(C))\equiv(B^{\prime},h(C^{\prime}))$, one also deduces that $B^{\prime\prime}$, $B$ are conditionally i.i.d. given $h(C^{\prime})$.
Furthermore, it follows from $(B^{\prime\prime}|A^{\prime},B^{\prime},h(C^{\prime}))\equiv(B^{\prime\prime}|h(C^{\prime}))$ that $(A^{\prime},B^{\prime})$ and $B^{\prime\prime}$ are conditionally independent given $h(C^{\prime})$.

### v2

@@ 7 Examples
In this section, we show details of some examples. Sometimes, we will write $\iota(x)=-x\log x$ for convenience.
Example 7.1.
Given $0\leq s\leq 1$, and given $n$ sufficiently large, we will construct i.i.d. random variables $X$ and $Y$ whose ranges are contained in $[0,1]$, such that $(X,Y)$ is marginally $(s,O(1);s,O(1))$-Frostman, and compare $H_{n}(X)$, $H_{n}(X+Y)$, $H_{n}(X^{2}+Y^{2})$ and $H_{n}(XY)$.
Indeed, let $m=m(n)\in\mathbb{N}$ and $l=l(n)\in\mathbb{N}\cup\{0\}$ be parameters to be chosen later. We suppose that $(X,Y)$ has the uniform distribution on $A\times A$, with
@@
Denote by $\mathcal{L}^{d}$ the $d$-dimensional Lesbegue measure. Then $\mathcal{L}^{1}(A)=2^{-l}$. We see that both $X$ and $Y$ have Note that $X$ has uniform distribution on $A$, i.e.,
@@
Claim 1. The random variable $(X,Y)$ is marginally $\left(s_{1},O(1);s_{2},O(1)\right)$-Frostman whenever $0\leq s_{1},s_{2}\leq\frac{m}{l+m}$.
@@
when $l/m$ tends to infinity. For $l=0$ and $s=1$, we assert by Claim 2 that $H_{n}(X)=n=ns$. Combining (12), (13) and Lemma 2.34, we conclude that
@@
$H(X^{\prime}+Y^{\prime})\leq\log\#\,range_{+}(X^{\prime}+Y^{\prime})=m+O(1),$
@@
Recall that $X^{\prime}=D_{m+l}(X)$ and $Y^{\prime}=D_{m+l}(Y)$. Since the function $(x,y)\mapsto x^{2}+y^{2}$ on $0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.18 that
@@
where $\mathcal{K}$ denotes the set consisting of integers $k$ such that $k=j_{1}^{2}+j_{2}^{2}$ for some $0\leq j_{1},j_{2}\leq 2^{m}-1$, and
@@
Since the function $(x,y)\mapsto xy$ on $0,1]^{2}$ is $O(1)$-Lipschitz, we obtain by Lemma 2.18 that
@@
where $\mathcal{K}^{\prime}$ denotes the set consisting of integers $k$ such that $k=j_{1}j_{2}$ for some $0\leq j_{1},j_{2}\leq 2^{m}-1$, and
@@
Example 7.2.
Let $X$ and $Y$ be the independent $(s,O(1))$-Frostman random variables defined in 7.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=\sqrt{X}$ and $\tilde{Y}=\sqrt{Y}$. By Example 7.1, one sees that
@@
Example 7.3.
Let $X$, $Y$ be the independent $(s,O(1))$-Frostman random variables defined in Example 7.1, with $n=m+l$ and $s=\frac{m}{l+m}$. Consider $\tilde{X}=1+X$, which is $(s,O(1))$-Frostman with range contained in $[1,2)$. Let $\tilde{Y}$ be an i.i.d. copy of $X$. Take $\phi(x,y)=(x+y)^{2}$, which is a degenerate quadratic form. By Example 7.1, one can deduce that
@@
and $Y$ has uniform distribution on $0,2^{-2m}]$. One can verify that $X$ is $(1/2,O(1))$-Frostman, and $Y$ is $(1/2,O(2^{m}))$-Frostman. We take $s=1/2$ and consider $n=2m$.
@@
Example 7.5.
This example is an variant of Example 7.4. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}^{2}$. Note that $X$ and $Y$ are dependent. The marginal distributions of $X$ and $Y$ are both the uniform distribution on $\bigcup\nolimits_{j=0}^{2^{m}-1}\big{[}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$. So both $X$ and $Y$ are $(1/2,O(1))$-Frostman. Let $W_{1}=(X_{1},Y_{1}),\,(X_{2},Y_{2})$ be i.i.d. copies of $(X,Y)$. We consider $n\leq 2m$.
@@
Moreover, for the random variable $W$, its Frostman constant $C$ can grow with $m$. If we consider $n=2m$ in Theorem 2.30, if $1<s<3/2$, then one has $H_{n}(|W_{1}-W_{2}|)=(1+o(1))n/2<nt$ for some suitably chosen $t$ with s/2<t<s1/2s/2<t<s-1/2. That is to say, the term $\log C$ in Theorem 2.30 is essential.
@@
Example 7.6.
This example is a further variant of Examples 7.4 and 7.5. Let $W=(X,Y)$, and suppose that the joint distribution of $(X,Y)$ is the uniform distribution on the set $A=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}\times\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$. Then the marginal distributions of $X$ and $Y$ are the uniform distribution on the sets $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}j2^{-m},\,j2^{-m}+2^{-2m}\big{)}$ and $A_{0}^{\prime}=\bigcup\nolimits_{j=0}^{2^{m}-1}\big{(}-j2^{-m},\,-j2^{-m}+2^{-2m}\big{)}$, respectively. So both $X$ and $Y$ are $(1/2,O(1))$-Frostman.
Claim 1. The random variable $(X,Y)$ is jointly $(1/4,1/4,O(1))$-Frostman, but is not conditionally $(s_{1},O(1);s_{2},O(2))$-Frostman for some $(s_{1},s_{2})\neq(0,0)$.
Claim 2. Taking $n=2m$, one has $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$, and $H_{n}(X,Y)=m+O(1)$.
In particular, if we apply Theorem 1.4 with $(X,Y)$, take $\alpha=-\beta$, $s_{1}=s_{2}=s=1/4$ and $n=2m$, then the second and third entropies in the maximum-term nearly vanish. So the first term $(H_{n}(X)+H_{n}(Y))/2$ is essential. In this example, it is $m+O(1)$, which is obviously larger than $ns=m/2$. Furthermore, we can not substitute $\frac{1}{2}H(X,Y)=m/2$ for $(H_{n}(X)+H_{n}(Y))/2$ as in Theorem 1.3.
Proof of Claim 1.
Note that $\mathcal{L}^{2}(A)=2^{-3m}$ and $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be a Borel set and $B(y,r)$ be any ball. When $0<r<2^{-2m}$, we have
$P(X\in E,Y\in B(y,r))\ll\mathcal{L}^{1}(E\cap A_{0})\cdot r\cdot 2^{3m},$
while
$P(X\in E)\leq\mathcal{L}^{1}\cdot 2^{m}.$
It follows that
$P(X\in E,Y\in B(y,r))\ll r^{1/2}P(X\in E).$
For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, one has
$P(X\in B(x,r_{1}),\,Y\in B(y,r_{2}))\ll\begin{cases}r_{1}r_{2}2^{3m},\quad&\text{if }0<r_{1}\leq r_{2}\leq 2^{-2m},\\ r_{1}2^{m},\quad&\text{if }0<r_{1}\leq 2^{-2m}<r_{2},\\ 2^{-m},\quad&\text{if }2^{-2m}<r_{1}\leq 2^{-m},\,r_{2}\geq r_{1},\\ r_{1},\quad&\text{if }2^{-m}<r_{1}\leq 1,\,r_{2}\geq r_{1},\\ 1,\quad&\text{if }r_{2}\geq r_{1}>1,\end{cases}$
and the other cases with $r_{2}<r_{1}$ can be obtained by symmetry. One can verify that all the terms on the right-hand side of above formula can be bounded by $r_{1}^{1/4}r_{2}^{1/4}$.
For the conditional Frostman condition, we take $E=B(x,2^{-2m})$ and $F=B(y,2^{-2m})$ such that $E\times F\subseteq A$. Then
$P(X\in E)=2^{-m}=P(X\in E,\,Y\in B(x,2^{-2m}))=2^{-m}.$
The only possible positive $s_{2}$ satisfying $2^{-m}\leq 2^{-m}(2^{-2m})^{s_{2}}$ is 0. Now it is not hard to see that $(X,Y)$ is $(s_{1},O(1);s_{2},O(1))$-Frostman only when $(s_{1},s_{2})=(0,0)$. ∎
Proof of Claim 2.
We use similar arguments as in the proof of Claim 1 in Example 7.5. Let $(X^{\prime},Y^{\prime})=D_{2m}(X,Y)$. Then $(X^{\prime},Y^{\prime})$ is the discrete uniform distribution on $\{(j2^{-m},-j2^{-m}:\,j=0,1,\ldots,2^{m}-1\}$. So $H_{n}(X^{\prime},Y^{\prime})=m$. Moreover, $range(X^{\prime}+Y^{\prime})=range(X^{\prime 2}-Y^{\prime 2})=\{0\}$, which implies that $H_{n}(X^{\prime}+Y^{\prime})=H_{n}(X^{\prime 2}-Y^{\prime 2})=0$. Applying similar relation as in (66), one asserts that $H_{n}(X+Y)=H_{n}(X^{2}-Y^{2})=O(1)$. Similarly, $H_{n}(X,Y)=m+O(1)$. ∎
Example 7.7.
Let
$B=\bigcup\limits_{i=0}^{2^{m}-1}\bigcup\limits_{j=0}^{2^{m}-1}[i2^{-m}+j2^{-2m},i2^{-m}+(j+1)2^{-2m})\times[j2^{-m}+i2^{-2m},i2^{-m}+(i+1)2^{-2m}).$
Then, for each given $x\in[0,1)$ (or $y\in[0,1)$), the set $\{y:\,(x,y)\in B\}$ (or $\{x:\,(x,y)\in B\}$, respectively) is exactly an interval of length $2^{-2m}$. Now Let
$A=\bigcup\limits_{i^{\prime},j^{\prime}=0}^{2^{m}-1}(i2^{-m},j2^{-m})+2^{-2m}\cdot B.$
Here $2^{-2m}\cdot B=\{(2^{-2m}x,2^{-2m}y):\,(x,y)\in B\}$. Suppose that $(X,Y)$ is a random variable that has uniform distribution on $A$.
Claim 1. The random variable $(X,Y)$ is marginally $(1/2,O(1);1/2,O(1))$-Frostman.
Claim 2. The random variable $(X,Y)$ is conditionally $(1/4,O(1);1/4,O(1))$-Frostman.
Claim 3. The random variable $(X,Y)$ is jointly $(1/2,1/2,O(1))$-Frostman.
Proof of Claim 1.
One can check that both $X$ and $Y$ have the uniform distribution on the set $A_{0}=\bigcup\nolimits_{j=0}^{2^{m}-1}\left[j2^{-m},\,j2^{-m}+2^{-m-l}\right)$. By Example 7.1, they are $(1/2,O(1))$-Frostman. ∎
Proof of Claim 2.
Note that $\mathcal{L}^{2}(A)=(2^{-4m})^{2}\cdot(2^{m})^{2}\cdot(2^{m})^{2}=2^{-4m}$. And $\mathcal{L}^{1}(A_{0})=2^{-m}$. Let $E$ be any Borel set, and $B(y,r)$ be any ball. Calculation reveals that
$P(X\in E)=\mathcal{L}^{1}(E\cap A_{0})\cdot 2^{m},$
and
$P(X\in E,y\in B(y,r))\ll\begin{cases}\mathcal{L}^{1}(E\cap A_{0})\cdot r\cdot 2^{4m},\quad&\text{if }0<r\leq 2^{-4m},\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}\cdot 2^{4m},\quad&\text{if }2^{-4m}<r\leq 2^{-m},\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}\frac{r}{2^{-m}}\cdot 2^{4m},\quad&\text{if }2^{-m}<r\leq 1,\\ \mathcal{L}^{1}(E\cap A_{0})\cdot 2^{-4m}2^{m}\cdot 2^{4m},\quad&\text{if }r>1.\end{cases}$
For all the cases, it satisfies that
$P(X\in E,y\in B(x,r))\ll r^{1/4}P(X\in E).$
Hence $Y$ is conditionally $(1/4,O(1))$-Frostman given $X$. Similarly, one can deduce that $X$ is conditionally $(1/4,O(1))$-Frostman given $Y$. Claim 2 then follows. ∎
Proof of Claim 3.
For any pair of balls $B(x,r_{1})$ and $B(y,r_{2})$, calculation reveal that
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))$
[…]
The rest cases with $r_{1}>r_{2}$ can be obtained by symmetry. In all cases, we have
$P(X\in B(x,y_{1}),y\in B(y,r_{2}))\ll r_{1}^{1/2}r_{2}{1/2}.$
The proof is completed. ∎
Example 7.8.
@@
Let $X,Y$ be the random variables defined in Example 7.1, and $X^{\prime}=D_{m+l}(X)$, $Y^{\prime}=D_{m+l}(Y)$, where the parameters are chosen as the following: for any given number $\eta>0$, let $m$ be any sufficiently large natural number, $l$ be a natural number with $(1+\eta/2)m\leq l\leq(1+\eta)m$, and $n=m+l$. By Example 7.1, the random variables $X$ and $Y$ are $(s,O(1))$-Frostman, where $s=\frac{m}{l+m}\in[\frac{1}{2+\eta},1/2)$. Also, the equalities in (65) hold.
@@
Write $i=i_{1}-i_{2}$ and $i^{\prime}=i_{1}^{\prime}-i_{2}^{\prime}$. Then $|i|,|i^{\prime}|<2^{m}$ and $|i+\sqrt{2}i^{\prime}|<2^{-(1+\eta/2)m}$. Combining the fact that $(i,i^{\prime})\neq(0,0)$, one can deduce that neither $i$ nor $i^{\prime}$ can be 0.
@@
Appendix A Appendix - Existence of Certain Random Variable
@@
Lemma A.1.
Let $(X,Y,Z)$ be a given random variable. Let $h$ be a function such that $h(Z)$ is a well-defined discrete random variable. Then there exists a random variable $(X^{\prime},Y^{\prime},Z^{\prime},Y^{\prime\prime})$ such that $(X^{\prime},Y^{\prime},Z^{\prime})\equiv(X,Y,Z)$, $(Y^{\prime\prime},h(Z^{\prime}))\equiv(Y,h(Z))$, and $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$. Moreover, $Y^{\prime\prime}$ and $Y^{\prime}$ are conditionally i.i.d. given $h(Z^{\prime})$. Furthermore, $(X^{\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $h(Z^{\prime})$.
@@
Suppose that $X,Y,Z$ are defined as measurable functions from a probability space $(\Omega,\mathcal{B},P)$ to some other measurable space. Now we construct a new probability space $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$. Let
$\Omega=\bigcup_{z\in range(h(Z))}\Omega_{z},\quad\Omega_{z}:=\{h(Z)=z\}=\{\omega\in\Omega:\,h(Z(\omega))=z\}.$
@@
$\Omega^{\prime}=\bigcup_{z}(\Omega_{z}\times\Omega_{z}).$
Then $(\omega,\nu)\in\Omega^{\prime}$ if and only if $\omega,\nu\in\Omega_{z}$ for some $z$, if and only if $h(Z(\omega))=h(Z(\nu))$. The $\sigma$-algebra $\mathcal{B}^{\prime}$ can be generated by the related Cartisian products of Borel sets in $\mathcal{B}$, and the probability measure $P^{\prime}$ can be induced by the following: if $E\times F$ is a Borel subset of $\Omega_{z}\times\Omega_{z}$ for some $z$, then
$P^{\prime}(E\times F)=\begin{cases}\frac{P(E)P(F)}{P(\Omega_{z})},\quad&\text{if }P(\Omega_{z})>0,\\ 0,&\text{if }P(\Omega_{z})=0.\end{cases}$
In the following, we assume without loss of generality that $P(\Omega_{z})>0$ for all $z\in range(h(Z))$. Now for any random variable $W$ defined on $(\Omega,\mathcal{B},P)$, let $W^{\prime}$ be the following random variable on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$:
$W^{\prime}(\omega,\nu)=W(\omega),\quad(\omega,\nu)\in\Omega^{\prime}.$
By this definition, one sees that $\big{(}f(W))^{\prime}=f(W^{\prime})$ for any admissible function $f$. In the following, we verify that $W^{\prime}\equiv W$. For any Borel set $L$,
$P^{\prime}(W^{\prime}\in L)$ $=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,W^{\prime}(\omega,\nu)\in L)$
$=\sum\limits_{z}P^{\prime}(\{(\omega,\nu)\in\Omega_{z}\times\Omega_{z}:\,W(\omega)\in L\})$
$=\sum\limits_{z}\frac{P(W\in L,h(Z)=z)P(h(Z)=z)}{P(h(Z)=z)}=P(W\in L).$
In particular, we can take $W=(X,Y,Z)$ to get $(X^{\prime},Y^{\prime},Z^{\prime})\equiv(X,Y,Z)$. All their corresponding marginal distributions are also the same. One has $h(Z)\equiv h(Z^{\prime})$ as well.
Next, we define $Y^{\prime\prime}$ as a function on $(\Omega^{\prime},\mathcal{B}^{\prime},P^{\prime})$, given by
$Y^{\prime\prime}(\omega,\nu)=Y(\nu),\quad(\omega,\nu)\in\Omega^{\prime}.$
@@
$P^{\prime}(Y^{\prime\prime}\in L_{1},W^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(\{(\omega,\nu)\in\Omega^{\prime}:\,Y^{\prime\prime}(\omega,\nu)\in L_{1},W^{\prime}(\omega,\nu)\in L_{2},h(Z^{\prime}(\omega,\nu))=z\})$
$\qquad=P^{\prime}(\{(\omega,\nu)\in\Omega_{z}\times\Omega_{z}:\,Y(\nu)\in L_{1},W(\omega)\in L_{2}\})=\frac{P(W\in L_{2},h(Z)=z)P(Y\in L_{1},h(Z)=z)}{P(h(Z)=z)}.$
When we take $W$ to be trivial (a constant function), then $P^{\prime}(Y^{\prime\prime}\in L_{1},h(Z^{\prime})=z)=P(B\in L_{1},h(Z)=z)$ and then $(Y^{\prime\prime},h(Z^{\prime}))\equiv(Y,h(Z))$. Moreover, the previous argument implies that $(W^{\prime},h(Z^{\prime}))\equiv(W,h(Z))$. We can calculate further to obtain
$P^{\prime}(Y^{\prime\prime}\in L_{1},W^{\prime}\in L_{2},h(Z^{\prime})=z)$ $=\frac{P(W\in L_{2},h(Z)=z)P(Y\in L_{1},h(Z)=z)}{P(h(Z)=z)}$
$=\frac{P^{\prime}(W^{\prime}\in L_{2},h(Z^{\prime})=z)P^{\prime}(Y^{\prime\prime}\in L_{1},h(Z^{\prime})=z)}{P^{\prime}(h(Z^{\prime})=z)}$
$=P^{\prime}(W^{\prime}\in L_{2},h(Z^{\prime})=z)P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z).$
So $P^{\prime}(Y^{\prime\prime}\in L_{1}|D^{\prime}\in L_{2},h(Z^{\prime})=z)=P^{\prime}(Y^{\prime\prime}\in L_{1}|h(Z^{\prime})=z)$, i.e. $(Y^{\prime\prime}|W^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$. Taking $W=(X,Y)$, one gets $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$.
Moreover, if we take $W=Y$, then $(Y^{\prime\prime}|Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$. That is to say, $Y^{\prime\prime}$ and $Y^{\prime}$ are conditionally independent given $h(Z^{\prime})$. Since $(Y^{\prime\prime},h(Z^{\prime}))\equiv(Y,h(Z))\equiv(Y^{\prime},h(Z^{\prime}))$, one also deduces that $Y^{\prime\prime}$, $B$ are conditionally i.i.d. given $h(Z^{\prime})$.
Furthermore, it follows from $(Y^{\prime\prime}|X^{\prime},Y^{\prime},h(Z^{\prime}))\equiv(Y^{\prime\prime}|h(Z^{\prime}))$ that $(X^{\prime},Y^{\prime})$ and $Y^{\prime\prime}$ are conditionally independent given $h(Z^{\prime})$.